OpenStudy (anonymous):

Within the lecture on session 1, it is stated that Y0 = 1/x0 (Y subzero One over x subzero). Why is that? I understood that Yo = f(1/x0). Also, in the example f(x) = 1/x, seems a statement rather that a truth.

OpenStudy (amistre64):

sub notations bring recurrsive things to mind

OpenStudy (amistre64):

not all of us know what lecture on session 1 refers to tho ....

OpenStudy (anonymous):

Sorry, folks. It is Calculus 1, First Class. Around minute 17:32. The statement is Yo = 1/xo. It is read, Y subzero = 1/X subzero). The question is why is this true?